help me
solve this equation
x+2y=3
x-y=4

Answer by Alan3354(69443)

solve this equation
x+2y=3
x-y=4
==============
ooooh, there are 2 equations.
If you want to solve the system:
----
x+2y=3
x -y=4
--------------------------Subtract